# UniGraph: Learning a Cross-Domain Graph Foundation Model From Natural Language

### v1

@@ 4.2 Graph Siamese Masked Autoencoders on TAGs
Enabling in-context learning. A dataset after generating embeddings, comprising node/edge/graph embedding-label pairs $({\bm{h}},y)$. An $N$-way $K$-shot transfer task involves a support set $\mathcal{S}=\{({\bm{h}}_{i},y_{i})\}_{i=1}^{N\times K}$ and a query set $\mathcal{Q}$. For each class $c$, the model computes an average embedding $\overline{{\bm{e}}_{c}}=\frac{1}{K}\sum_{({\bm{h}},y)\in\mathcal{S},y=c}{\bm{h}}$. The classification of a query data sample ${\bm{h}}_{q}$ in $\mathcal{Q}$ is determined by comparing its embedding to these average embeddings using the cosine similarity. The performance of the model is evaluated using a metric $\mathcal{M}$, which could be accuracy, defined as $\mathcal{M}=\frac{1}{|\mathcal{Q}|}\sum_{{\bm{h}}_{q}\in\mathcal{Q}}\mathbbm{1}\left[\arg\max_{c\in\{1,\ldots,N\}}\frac{{\bm{h}}_{q}\cdot\overline{e_{c}}}{\|{\bm{h}}_{q}\|\|\overline{e_{c}}\|}=y_{q}\right]$.
4.3 Graph Instruction Tuning
Graphs from different domains have distinct label spaces, making it challenging to directly transfer to unseen graphs without fine-tuning the pre-trained graph models using a substantial number of labels (Hou et al., 2023). Inspired by Vision-Language Models (VLMs) (Zhu et al., 2023), to enhance zero-shot capabilities, we propose using a graph instruction tuned open-ended generative LLM to unify the label spaces of different graphs. In instruction tuning, we provide textual instructions to the model which vary for different domains, to facilitate adaptation to each domain. To train the model to use these instructions, we fine-tune the model on instruction tuning datasets where we have labels.
Instruction prompts. For instruction tuning, we design prompt templates that include graph embeddings, graph structural information, and natural language instructions, as summarized in Appendix G.
Training. Our instruction tuning pipeline is shown in Figure 1. Given a query graph, we first generate embeddings for it using our pretrained graph model $f_{\theta}$. We apply a linear projector to the graph embeddings to map them into the LLM’s embedding space. Then we combine graph embeddings and natural language instructions as inputs to the LLM. We select Llama-7B (Touvron et al., 2023a) as our LLM and adopt LoRA (Hu et al., 2021a) for fine-tuning the LLM while keeping the word embedding frozen. We fine-tune the LLM only to generate the label in natural language and the loss is computed only on the predicted target.
Table 3: Experiment results in few-shot transfer. We report accuracy (%) for both datasets. Performance of OFA, which is directly trained on test datasets with labels, is marked in gray. Prodigy is pre-trained on the MAG240M or Wiki. UniGraph and other self-supervised learning baselines are pre-trained on ogbn-Papers100M, and then evaluated on individual target dataset.
@@ 4.3 Graph Instruction Tuning
Table 4: Experiment results in zero-shot transfer. We report accuracy (%) for all datasets. UniGraph-IT represents our self-supervised pre-trained model with our instruction tuned LLM.
[…]
@@
Table 5: Ablation studies of UniGraph key components.
@@
Table 6: Ablation studies of UniGraph pre-training datasets. ’citation’, ’KG’, and ’MOL’ respectively refer to ogbn-Papers100M, FB15K237, and ChEMBL.
@@
Table 7: Analysis of LMs and GNNs choices.
@@
Table 8: Efficiency comparison of our method versus only using LM for pretraining on an NVIDIA A100 (40G) GPU.
@@
5 Experiments
5.1 Self-Supervised Representation Learning
Setup. We adopt the most commonly used linear probing protocol to evaluate the representation learning ability of self-supervised pre-trained models on unseen datasets. We train a linear classifier at the top the embedding from frozen model. Our model and all self-supervised learning baseline methods are first pre-trained on the large-scale citation network ogbn-Papers100M (Hu et al., 2020). Then, we evaluate on nine graphs from five domains with different tasks. Detailed settings and hyper-parameters can be found in Appendix C.
For the baselines, we compare UniGraph with state-of-the-art generative graph self-supervised learning methods: GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023), constrastive methods: DGI (Veličković et al., 2018b) and BGRL (Thakoor et al., 2021). As they are not designed for cross-domain purposes, we utilize shallow LM word2vec (Mikolov et al., 2013) and pre-trained LM DeBERTa-base (He et al., 2020) to unify the input node features of different graphs. Consistent with our approach, all baseline methods use GAT (Veličković et al., 2018a) as the backbone GNN. We also compare with supervised models GCN (Kipf & Welling, 2016a) and GAT, which are separately trained on the target datasets. For baseline methods that take TAGs as input, we select GIANT-XRT, which fine-tunes language models using graph structure, and GIANT-XRT+GraphMAE2, which conducts second pre-training based on embeddings generated by GIANT-XRT.
Results. Table 2 presents the results. We interpret these results from three perspectives: (1) UniGraph outperforms state-of-the-art graph self-supervised learning methods by a large margin. This indicates that our framework possesses a stronger generalization ability in cross-domain graph learning scenarios, enabling it to generate more discriminative embeddings for unseen graphs. (2) Compared to using pre-processed features, learning directly from TAGs is more advantageous for cross-domain transfer. UniGraph and GIANT-XRT, which take text as input, demonstrate stronger performance than GNN-based self-supervised learning methods that use LMs to pre-encode text features. (3) As a single pre-trained model applied to various downstream datasets, UniGraph exhibits better or comparable performance than supervised learning models trained directly on those downstream datasets. This further suggests the feasibility and effectiveness of training a self-supervised graph foundation model.
5.2 Few-Shot Transfer
Setup. In this part, we evaluate the ability of the pre-trained models to perform few-shot in-context transfer without updating the model parameters. For baseline methods, in addition to the pre-trained models mentioned in section 5.1, we also compared two latest graph in-context learning methods: the self-supervised pre-training method Prodigy (Huang et al., 2023) and the supervised pre-training method OFA (Liu et al., 2023a). They each utilize different LMs to unify the input features of different graphs. Unlike our setting, Prodigy is pre-trained on the MAG240M (Hu et al., 2021b) or Wiki datasets (Huang et al., 2023) for corresponding downstream Arxiv or FB15K237 datasets, while OFA is pre-trained on Arxiv, FB15K237, and ChEMBL.
For evaluation, we strictly follow the setting of Prodigy (Huang et al., 2023). For an N-way K-shot task, we adopt the original train/validation/test splits in each downstream classification dataset, and construct a $K$-shot prompt for test nodes (or edges) from the test split by randomly selecting $K$ examples per way from the train split. By default in all experiments, we sample 500 test tasks with 3-shot prompts. Detailed settings can be found in Appendix C.
Results. In table 3, the results demonstrate that our UniGraph framework consistently outperforms all the baselines. In particular, compared to Prodigy and OFA, which are pre-trained on the same tasks as the downstream tasks, our model still achieves superior performance. This suggests that our graph foundation model is capable of learning effective general knowledge from pre-training tasks and can learn in the context of downstream tasks without tuning the model parameters.
5.3 Zero-Shot Transfer
Setup. For zero-shot transfer, we mainly compare with general-purpose LLMs such as Llama (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) and vicuna-7B-v1.5 (Chiang et al., 2023), as well as graph-based LLM methods GraphGPT (Tang et al., 2023). Also, we compare with OFA mentioned in section 5.2. Consistent with OFA, we select ogbn-Arxiv and FB15K237 as instruction tuning datasets. Note that there is no overlap between their label categories and those of the downstream datasets. We report the average accuracy on the original test set with 3 random initializations for each downstream dataset.
Results. In table 4, we can observe that our proposed UniGraph significantly outperforms open-source LLMs, confirming that our framework can effectively align graph embeddings with natural language representations. Also, LLMs are capable of learning transferable graph structure and graph learning task knowledge from our designed graph instruction tuning.
5.4 Model Analysis
@@
Ablation on pre-training datasets. How to build pre-training datasets has been a key research topic in training foundational models (Bommasani et al., 2021). Table 6 shows the impact of including graphs from different domains in the pre-training datasets on downstream performance. We can observe that pre-training on graphs from the same domain enhances the performance of downstream tasks. This suggests that in-domain transfer remains simpler than cross-domain transfer. However, this paper primarily focuses on the generalization ability of the graph foundation model in cross-domain transfer. And for a fair comparison with baselines that train only on a single dataset, we restrict our pre-training to citation network ogbn-Papers100M. We leave the construction of diverse and high-quality pre-training datasets for future research.
@@
Efficiency analysis. In Table 8, we can observe that the time and space overhead of training our method is comparable to only training an LM with the MLM task. We can conclude that the computation cost of our framework is dominated by the LM cost, and our method’s running time is similar to other LM-based approaches if they use similar LMs. More efficiency analysis can be found in Appendix A.
6 Conclusion
In this work, we present UniGraph framework, aimed at designing and training a novel graph foundation model that enables cross-domain generalization. We have successfully demonstrated that through the integration of LMs and GNNs, our framework can effectively learn and transfer knowledge across diverse graph domains. The experimental results, covering a wide range of graph learning tasks and scenarios, validate the robustness and versatility of UniGraph. This work not only addresses a critical gap in graph learning but also lays down a foundational framework that can be further explored and refined for broader applications.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Efficiency Analysis.
@@ 7 Impact Statements
Time Complexity. The LM processes sequences with a complexity that scales with the square of the sequence length and linearly with the dimensionality of the embeddings. For a given sequence, the complexity is $O(L^{2}d+Ld^{2})$, where $L$ is the sequence length and $d$ is the dimensionality of the feature vectors. Processing $N$ such sequences within a batch, the time complexity becomes $O(N\cdot(L^{2}d+Ld^{2}))$.
The GNN’s time complexity primarily depends on the feature dimensionality and the structure of the graph. For simplicity, assuming the complexity of aggregating features for each node is proportional to the dimensionality of the features and the degree of connectivity, the GNN encoder’s complexity can be approximated as $O(N\cdot d^{2})$ if we consider only the immediate neighborhood aggregation. For dense graphs or more complex aggregation schemes, this might scale further as $O(N\cdot d^{2}+N^{2}\cdot d)$, but it remains secondary to the LM’s complexity in most configurations.
For latent loss calculation, the complexity is similar to the additional cost for the cosine similarity computation, which is $O(N\cdot d)$.
The dominating factor in the time complexity is the processing by the LM, leading to $O(N\cdot(L^{2}d+Ld^{2}))$, with the GNN adding a lesser, though not negligible, additional cost.
Space Complexity. The space complexity for the LM includes the storage for intermediate activations and the model parameters. For activations, the primary concern is the size of the sequence by the feature dimensionality, $O(N\cdot L\cdot d)$, and the self-attention mechanism’s quadratic dependence on sequence length, $O(N\cdot L^{2})$. Thus, the space complexity for the LM part is $O(N\cdot L^{2}+N\cdot L\cdot d)$.
The space requirement for the GNN is lower, mainly involving the storage of node features before and after processing. This can be approximated as $O(N^{2}+d^{2}+N\cdot d)$ for dense graphs, assuming each node’s features are stored in a vector of dimension $d$.
The overall space complexity is mainly influenced by the LM’s requirements for storing activations and parameters, with $O(N\cdot L^{2}+N\cdot L\cdot d)$ being the primary factor. The GNN’s space requirements add minimally to the total.
Appendix B Datasets

### v2

@@ 4.2. Graph Siamese Masked Autoencoders on TAGs
Enabling in-context learning. A dataset after generating embeddings, comprising node/edge/graph embedding-label pairs $({\bm{h}},y)$. An $N$-way $K$-shot transfer task involves a support set $\mathcal{S}=\{({\bm{h}}_{i},y_{i})\}_{i=1}^{N\times K}$ and a query set $\mathcal{Q}$. For each class $c$, the model computes an average embedding $\overline{{\bm{e}}_{c}}=\frac{1}{K}\sum_{({\bm{h}},y)\in\mathcal{S},y=c}{\bm{h}}$. The classification of a query data sample ${\bm{h}}_{q}$ in $\mathcal{Q}$ is determined by comparing its embedding to these average embeddings using the cosine similarity. The performance of the model is evaluated using a metric $\mathcal{M}$, which could be accuracy, defined as:
(7) $\mathcal{M}=\frac{1}{|\mathcal{Q}|}\sum_{{\bm{h}}_{q}\in\mathcal{Q}}\mathbbm{1}\left[\arg\max_{c\in\{1,\ldots,N\}}\frac{{\bm{h}}_{q}\cdot\overline{e_{c}}}{\|{\bm{h}}_{q}\|\|\overline{e_{c}}\|}=y_{q}\right].$
Table 2. Experiment results in few-shot transfer. We report accuracy (%) for both datasets. Performance of OFA, which is directly trained on test datasets with labels, is marked in gray. Prodigy is pre-trained on the MAG240M or Wiki. UniGraph and other self-supervised learning baselines are pre-trained on ogbn-Papers100M, and then evaluated on individual target dataset.
@@
4.3. Graph Instruction Tuning
Graphs from different domains have distinct label spaces, making it challenging to directly transfer to unseen graphs without fine-tuning the pre-trained graph models using a substantial number of labels (Hou et al., 2023). To enhance zero-shot capabilities, we propose using a graph instruction tuned open-ended generative LLM to unify the label spaces of different graphs. In instruction tuning, we provide textual instructions to the model which vary for different domains, to facilitate adaptation to each domain. To train the model to use these instructions, we fine-tune the model on instruction tuning datasets where we have labels.
Instruction prompts. For instruction tuning, we design prompt templates that include graph embeddings, graph structure, and natural language instructions, as summarized in Appendix G.
Training. Our instruction tuning pipeline is shown in Figure 1. Given a query graph, we first generate embeddings for it using our pretrained model $f_{\theta}$. We apply a linear projector to map the graph embeddings into the LLM’s embedding space. Then we combine graph embeddings and natural language instructions as inputs to the LLM. We select Llama-7B (Touvron et al., 2023a) as our LLM and adopt LoRA (Hu et al., 2021b) for fine-tuning the LLM while keeping the word embeddings frozen. We fine-tune the LLM to generate labels in natural language and the loss is computed only on the predicted target.
Table 3. Experiment results in zero-shot transfer. We report accuracy (%) for all datasets. UniGraph-IT represents our self-supervised pre-trained model with our instruction tuned LLM.
Cora PubMed Products1 Wiki-CS WN18RR
@@ 4.3. Graph Instruction Tuning
  • 2
    Due to the difference between the pre-trained datasets used in their original paper and ours, we re-run the experiments using their open-source code. Comparisons between UniGraph and their method using their original pre-trained datasets can be found in Appendix J.
Table 4. Performance comparison with dataset-specific graph self-supervised learning.
Method Cora PubMed Arxiv Products Papers100M Wiki-CS FB15K237 WN18RR HIV PCBA ChEMBL
DGI-single 69.24$\pm$0.42 68.24$\pm$0.24 69.76$\pm$0.33 78.11$\pm$0.24 55.21$\pm$0.49 77.76$\pm$0.24 85.27$\pm$0.64 52.04$\pm$0.22 68.42$\pm$0.90 51.24$\pm$0.87 61.24$\pm$0.53
BGRL-single 71.24$\pm$0.11 69.01$\pm$0.34 70.87$\pm$0.24 79.01$\pm$0.32 64.45$\pm$0.32 78.24$\pm$0.24 87.24$\pm$0.25 72.11$\pm$0.68 68.23$\pm$0.33 55.82$\pm$1.23 62.04$\pm$0.54
GraphMAE-single 73.24$\pm$0.76 69.23$\pm$0.97 72.01$\pm$0.24 78.24$\pm$0.37 65.43$\pm$0.76 79.24$\pm$0.22 86.35$\pm$0.35 72.34$\pm$0.24 69.23$\pm$0.23 56.22$\pm$1.87 63.28$\pm$0.76
GraphMAE2-single 72.23$\pm$0.33 69.87$\pm$0.43 71.21$\pm$0.24 79.11$\pm$0.24 66.24$\pm$0.54 79.21$\pm$0.34 87.22$\pm$0.24 70.45$\pm$0.24 68.34$\pm$0.34 56.87$\pm$0.43 63.54$\pm$0.73
UniGraph-cross 81.43$\pm$0.55 74.33$\pm$0.23 72.91$\pm$0.42 80.11$\pm$0.23 67.89$\pm$0.21 79.98$\pm$1.21 94.81$\pm$1.32 85.45$\pm$0.34 71.23$\pm$1.93 57.67$\pm$0.85 64.29$\pm$1.01
UniGraph-single 84.23$\pm$0.24 80.11$\pm$0.21 73.97$\pm$0.22 82.24$\pm$0.24 67.89$\pm$0.21 81.22$\pm$0.24 95.24$\pm$0.23 87.21$\pm$0.76 75.24$\pm$1.24 60.23$\pm$0.11 65.32$\pm$0.21
5. Experiments
5.1. Self-Supervised Representation Learning
Setup. We adopt the most commonly used linear probing protocol to evaluate the representation learning ability of self-supervised pre-trained models on unseen datasets. We train a linear classifier on top of the embeddings from a frozen model. Our model and all self-supervised learning baseline methods are first pre-trained on the large-scale citation network ogbn-Papers100M (Hu et al., 2020). Then, we evaluate on nine graphs from five domains with different tasks. Detailed settings and hyper-parameters can be found in Appendix C.
For the baselines, we compare UniGraph with state-of-the-art generative graph self-supervised learning methods: GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023), constrastive methods: DGI (Veličković et al., 2018b) and BGRL (Thakoor et al., 2021). As they are not designed for cross-domain purposes, we utilize shallow LM word2vec (Mikolov et al., 2013) and pre-trained LM DeBERTa-base (He et al., 2020) to unify the input node features of different graphs. Consistent with our approach, all baseline methods use GAT (Veličković et al., 2018a) as the backbone GNN. We also compare with supervised models GCN (Kipf and Welling, 2016a) and GAT, which are separately trained on the target datasets. For baseline methods that take TAGs as input, we select GIANT-XRT, which fine-tunes language models using graph structure, and GIANT-XRT+GraphMAE2, which conducts second pre-training based on embeddings generated by GIANT-XRT.
Results. Table 1 presents the results. We interpret these results from three perspectives: (1) UniGraph outperforms state-of-the-art graph self-supervised learning methods by a large margin. This indicates that our framework possesses a stronger generalization ability in cross-domain graph learning scenarios, enabling it to generate more discriminative embeddings for unseen graphs. (2) Compared to using pre-processed features, learning directly from TAGs is more advantageous for cross-domain transfer. UniGraph and GIANT-XRT, which take text as input, demonstrate stronger performance than GNN-based methods that use LMs to pre-encode text features. (3) As a single pre-trained model applied to various downstream datasets, UniGraph exhibits better or comparable performance than supervised learning models trained directly on those downstream datasets. This further suggests the feasibility and effectiveness of training a self-supervised graph foundation model.
5.2. Few-Shot Transfer
Setup. In this part, we evaluate the ability of the pre-trained models to perform few-shot in-context transfer without updating the model parameters. For baseline methods, in addition to the pre-trained models mentioned in section 5.1, we also compared two latest graph in-context learning methods: the self-supervised pre-training method Prodigy (Huang et al., 2023) and the supervised pre-training method OFA (Liu et al., 2023a). They each utilize different LMs to unify the input features of different graphs. Unlike our setting, Prodigy is pre-trained on the MAG240M (Hu et al., 2021a) or Wiki datasets (Huang et al., 2023) for corresponding downstream Arxiv or FB15K237 datasets, while OFA is pre-trained on Arxiv, FB15K237, and ChEMBL. We also compare UniGraph with graph prompt learning methods, such as GPF (Fang et al., 2024), All in One (Sun et al., 2023), and GraphPrompt (Liu et al., 2023b). Additionally, we compare UniGraph with graph meta-learning methods, including GPN (Ding et al., 2020), TENT (Wang et al., 2022b), and GLITTER (Wang et al., 2022a).
For evaluation, we strictly follow the setting of Prodigy (Huang et al., 2023). For an N-way K-shot task, we adopt the original train/validation/test splits in each downstream classification dataset, and construct a $K$-shot prompt for test nodes (or edges) from the test split by randomly selecting $K$ examples per way from the train split. By default in all experiments, we sample 500 test tasks with 3-shot prompts.
Results. In table 2, the results demonstrate that our UniGraph framework consistently outperforms all the baselines. In particular, compared to Prodigy and OFA, which are pre-trained on the same tasks as the downstream tasks, our model still achieves superior performance. This suggests that our graph foundation model is capable of learning effective general knowledge from pre-training tasks and can learn in the context of downstream tasks. For graph prompt learning and graph meta learning methods, we can observe that although these baseline methods are fine-tuned on downstream datasets, UniGraph is still able to outperform these baselines without modifying the model parameters.
5.3. Zero-Shot Transfer
Setup. For zero-shot transfer, we mainly compare with general-purpose LLMs such as Llama (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) and vicuna-7B-v1.5 (Chiang et al., 2023), as well as graph-based LLM methods [^1] Due to the use of different pre-training datasets, comparisons with other graph zero-shot learning baselines, such as LLaGA (Chen et al., 2024) and OpenGraph (Xia et al., 2024), can be found in Appendix J. GraphGPT (Tang et al., 2023), ZeroG (Li et al., 2024) and G2P2 (Wen and Fang, 2023). Also, we compare with OFA mentioned in section 5.2. Consistent with OFA, we select ogbn-Arxiv and FB15K237 as instruction tuning datasets. Note that there is no overlap between their label categories and those of the downstream datasets. We report the average accuracy on the original test set with 3 random initializations for each downstream dataset.
Results. In table 3, we can observe that our proposed UniGraph significantly outperforms open-source LLMs, confirming that our framework can effectively align graph embeddings with natural language representations. Also, LLMs are capable of learning transferable graph structure and graph learning task knowledge from our designed graph instruction tuning. Additionally, compared to other graph zero-shot learning methods, UniGraph also demonstrates consistent advantages.
5.4. Comparisons With Dataset-Specific Graph Self-Supervised Learning
Table 4 shows the performance under self-supervised representation learning setting, where ”-single” indicates self-supervised training on the target dataset, and ”-cross” signifies self-supervised pre-training on the pre-training dataset Papers100M. For DGI, BGRL, GraphMAE, and GraphMAE2, word2vec features are used as model inputs, whereas for UniGraph, text is directly used as input.
From the experimental results, we observe that UniGraph still outperforms dataset-specific Graph SSL, further illustrating the potential and feasibility of foundation models. At the same time, the performance of UniGraph with dataset-specific self-supervised training exceeds that of cross-domain UniGraph to some extent. This indicates the effectiveness of the proposed SSL algorithm and suggests that there is still room for improvement in cross-domain pre-training. Note that cross-domain UniGraph avoids the need for collecting additional task-specific data for training.
Table 5. Comparison of computational costs and performance on ogbn-Arxiv and ogbn-Papers100M.
Dataset Method Pre-training Time Downstream Training Time Downstream Inference Time Test Accuracy
ogbn-Arxiv (169,343 nodes) GAT - 23.2 mins 5.5 mins 70.89 $\pm$ 0.43
GraphMAE2 - 4.8 h 5.1 mins 71.21 $\pm$ 0.24
UniGraph 23.4 h - 10.2 mins 72.91 $\pm$ 0.42
ogbn-Papers100M (111,059,956 nodes) GAT - 6.8 h 23.1 mins 65.98 $\pm$ 0.23
GraphMAE2 - 20.1 h 24.3 mins 66.24 $\pm$ 0.54
UniGraph 23.4 h - 41.2 mins 67.89 $\pm$ 0.21
Table 6. Ablation studies of UniGraph key components.
@@ 5.4. Comparisons With Dataset-Specific Graph Self-Supervised Learning
Table 7. Ablation studies of UniGraph pre-training datasets. ’citation’, ’KG’, and ’MOL’ respectively refer to ogbn-Papers100M, FB15K237, and ChEMBL.
@@
Table 8. Analysis of LMs and GNNs choices.
@@
Table 9. Efficiency comparison of our method versus only using LM for pretraining on an NVIDIA A100 (40G) GPU.
@@
5.5. Model Analysis
@@
Ablation on pre-training datasets. Table 7 shows the impact of including graphs from different domains in the pre-training datasets on downstream performance. We can observe that pre-training on graphs from the same domain enhances the performance of downstream tasks. This suggests that in-domain transfer remains simpler than cross-domain transfer. However, this paper primarily focuses on the generalization ability of the graph foundation model in cross-domain transfer. For fair comparison with baselines that train only on a single dataset, we restrict our pre-training to ogbn-Papers100M.
@@
5.6. Efficiency Analysis.
@@
Time complexity. The time complexity for pre-training is dominated by the LM processing, which scales as $O(N\cdot(L^{2}d+Ld^{2}))$, where $d$ is the dimensionality of the embeddings. The GNN adds a complexity of $O(N\cdot d^{2})$ for immediate neighborhood aggregation, potentially increasing to $O(N\cdot d^{2}+N^{2}\cdot d)$ for dense graphs. Latent loss calculation, such as cosine similarity, adds $O(N\cdot d)$. Thus, the overall complexity is primarily driven by the LM with $O(N\cdot(L^{2}d+Ld^{2}))$, with the GNN contributing a secondary, though not negligible, cost.
Space complexity. The space complexity for the LM includes storage for intermediate activations and model parameters, amounting to $O(N\cdot L\cdot d)$ for activations and $O(N\cdot L^{2})$ for the self-attention mechanism, leading to a total of $O(N\cdot L^{2}+N\cdot L\cdot d)$. For the GNN, the space requirement is primarily for storing node features, approximated as $O(N^{2}+d^{2}+N\cdot d)$ for dense graphs. Therefore, the overall space complexity is dominated by the LM, with $O(N\cdot L^{2}+N\cdot L\cdot d)$, while the GNN adds a relatively smaller contribution.
Pretraining efficiency. In Table 9, we can observe that the time and space overhead of training UniGraph is comparable to only training an LM with the MLM task. We can conclude that the computation cost of our framework is dominated by LM cost, and our method’s running time is similar to other LM-based methods if they use similar LMs.
Computational costs as a foundation model. UniGraph, designed as a foundation model, incurs significant computational costs primarily during the pre-training phase. However, it offers the advantage of applicability to new datasets in the inference phase without necessitating retraining. We conduct comparisons of the training/inference costs between our model and GNN-based models. GAT (Veličković et al., 2018a) is a supervised trained GNN. GraphMAE2 (Hou et al., 2023) is a self-supervised learning method with GAT as the backbone network. We choose ogbn-Arxiv and ogbn-Papers100M, two datasets of different scales for experiments. From the results in the table 5, we can observe that although UniGraph has a long pre-training time, its inference time on downstream datasets is comparable/shorter than the training plus inference time of GNN-based methods. This advantage further increases with the size and potential quantity of downstream datasets. The same conclusion also applies to space complexity. Although LM has a larger number of parameters, since we only need to perform inference on the downstream dataset, we avoid the additional space occupation in the backward propagation during training.
6. Conclusion
In this work, we present UniGraph framework, aimed at designing and training a novel foundation model for TAGs that enables cross-domain generalization. We have demonstrated that through large-scale pre-training, our framework can effectively learn and transfer knowledge across diverse graph domains. The experimental results, covering a wide range of graph learning tasks and scenarios, validate the robustness and versatility of UniGraph. This work not only addresses a critical gap in graph learning but also lays down a foundational framework that can be further explored and refined for broader applications.
Appendix A Datasets